OpenStudy (anonymous):

What is the slope and y-intercept for the line represented by this equation? 5x+13y= -13

OpenStudy (saifoo.khan):

For slope, make "y" the subject and then the value for "x" is the slope. [PS y = mx + b is the equation for slope]

OpenStudy (anonymous):

The slope is given by the coefficient in front of x (the number in front of x) when the equation is in standard form y=mx+b. So solve for y and see what you get :-)

OpenStudy (saifoo.khan):

5x+13y= -13 13y = -13 - 5x DIvide 13 from both sides, \[ y = \frac{-5x - 13}{13}\]\[y = -\frac{-5}{13}x - \frac{13}{13}\]Now compare it with y = mx +b therefore, slope = -5/13

OpenStudy (anonymous):

:-/ spoiled the suspense for sergeantpepper (btw sergeantpepper "solve for y" means get your equation to be y= "something").

OpenStudy (anonymous):

is the y-intercept -13 or -1?

OpenStudy (anonymous):

the y-intercept is b in the equation y=mx+b (note that when x=0 we have y=b so that's when we go through the y-axis-when x=0, hence the name the y-intercept).

OpenStudy (anonymous):

so the y-intercept is -13?

OpenStudy (anonymous):

it is -13/13=1